OpenStudy (anonymous):

Simplify this expression. 12x2(3x4 - 2x) Select one of the options below as your answer: A. 36x6 - 2x B. 36x6 - 24x3 C. 36x8 - 2x D. 36x8 - 24x3

Parth (parthkohli):

Do you know how to distribute?

Parth (parthkohli):

\( \color{Black}{\Rightarrow (12x^2 \times 3x^4) - (12x^2 \times 2x) }\)

Parth (parthkohli):

I guess you can manage now.

OpenStudy (anonymous):

12x2(3x4 - 2x) B. 36x6 - 24x3

OpenStudy (maheshmeghwal9):

plz make me understand it:)any1

OpenStudy (maheshmeghwal9):

step by step plz:)

OpenStudy (maheshmeghwal9):

i am not getting it.

OpenStudy (anonymous):

12*3=?

OpenStudy (maheshmeghwal9):

36

OpenStudy (anonymous):

ok good x^2*x^4=

OpenStudy (maheshmeghwal9):

\[x^6\]

OpenStudy (anonymous):

u always add the x's when ur multiplying

Parth (parthkohli):

\( \color{Black}{\Rightarrow 12x^2 \times 3x^4 = 12 \times 3 \times x \times x\times x\times x \times x\times x = 36x^6 }\) Think of it like this :)

OpenStudy (maheshmeghwal9):

yes!@ahaines14

OpenStudy (maheshmeghwal9):

Oh thanx a lot for helping me:) @ahaines14 & @ParthKohli

OpenStudy (anonymous):

ur welcome hun:)

OpenStudy (maheshmeghwal9):

Thank u!